A car number plate consists of a letter, followed by five numbers, and ended with three letters. How many car number plates can be formed? if required there must not be the same letter and no equal lift, and how many car number plates can be formed ?
1
Expert's answer
2020-05-04T19:58:44-0400

There are 26 choices for the letter. Also there are 10 choices for number, so we have: N=26*105*263=45697600000

And if required that all letters must be different - then 26 choices decreased each time when we use letter, so N=26*105*25*24*23= 35880000000